# Estimating excess mortality during the Covid-19 pandemic in Aotearoa New Zealand

### v1

@@ 1 Introduction
Aotearoa New Zealand used a combination of border and community control measures to minimise transmission of SARS-CoV-2 until high vaccine coverage was achieved in late 2021 [1, 2]. Following establishment of the B.1.1.529 (Omicron) variant in January 2022 [3], New Zealand experienced a series of Covid-19 waves [4]. By 31 December 2023, there had been a total of 750 Covid-19-attributed deaths per million people [5], one of the lowest rates in the world.
The number of Covid-19-attributed deaths is one measure of the pandemic’s impact on mortality, but has limitations. It could be influenced by factors such as test availability, how Covid-19-attributed deaths are distinguished from incidental deaths in Covid-19 positive individuals, and processes for recording multiple causes of death. Also Covid-19-attributed deaths only measure the direct impact of the disease on mortality. This ignores indirect effects, such as deaths that occurred or were prevented as a result of pandemic control measures, and deaths that would have occurred due to another cause within a given time period.
@@
The Karlinsky and Kobak [7] baseline relies on extrapolating the pre-pandemic trend in raw death counts. This trend may, in some cases, mainly reflect the trend in total population size if the crude mortality rate (i.e. deaths per capita) does not vary greatly. Gibson [12] correctly pointed out that, in such cases, if the trend in total population size abruptly changes, then the Karlinsky and Kobak baseline will be systematically biased. There was a sharp drop in the rate of population growth in New Zealand in 2020 due to reduced international travel. Gibson [12] argued that controlling for population size by fitting to the trend in crude mortality rate, rather than raw deaths, led to a lower baseline than that of Karlinsky and Kobak [7] and, therefore, a higher estimate of excess mortality.
However, the trend in raw deaths may be also be influenced by population ageing, which would generally lead to an upward trend in crude mortality rate. For this reason, where age-stratified data are available, it is preferable to control for age when estimating mortality baselines [13]. This will limit international comparisons to jurisdictions where age-stratified data are available, but will provide more accurate estimates. Neither the method of Gibson [12] nor Kung et al. [11] controlled for age. Reliable estimates of pandemic excess mortality in New Zealand that control for the combined effects of population growth and ageing are lacking.
@@
Both methods account for the effects of changes in population size and age structure, which means they provide more reliable estimates of excess mortality than methods that do not do this. We use the quasi-Poisson regression model as our primary method as this allows us to disaggregate results to monthly time periods and by age and sex. We compare aggregated yearly results from this method with the standardised mortality rate linear regression to benchmark our results against a simpler model. We then compare the estimated excess to the number of Covid-19-attributed deaths over time and in different subgroups. This enables us to estimate whether there was any net non-Covid-19 excess mortality during this period.
@@ 2 Methods
We used two methods for constructing a mortality baseline from pre-pandemic data: (1) a quasi-Poisson regression model for monthly death counts including age, sex and month of the year as predictors, based closely on the model of the UK’s Office for National Statistics [15]; (2) a linear regression fitted to the age- and sex-standardised annual mortality rate. See Supplementary Material Sec. 1.2 for full model specification.
@@ 3 Results
Figure 1: Total monthly all-cause deaths from January 2010 to December 2023 (blue points) along with the quasi-Poisson regression model fitted to data on month deaths from January 2014 to December 2019. Vertical dashed lines show the fitting period. Solid blue curve shows the mean monthly deaths according to the fitted model; shaded band shows the 95% CI.
In the pre-pandemic period, all-cause monthly deaths had a pronounced seasonal pattern, with relatively high deaths during the winter (i.e. June–August) respiratory disease season (Figure 1). The quasi-Poisson regression model captured the seasonality and the gradually increasing trend in deaths. The age- and sex-standardised yearly all-cause mortality rate varied between 7.0 and 7.9 per 1000 between 2010 and 2019, with a decreasing trend (Figure 2). The modelled standardised yearly mortality rate was very similar under the two methods.
Figure 2: Age- and sex-standardised yearly all-cause mortality rate per 1000 people (open circles), with the results of the quasi-Poisson regression (QPR) model (solid blue curve = mean, shaded blue band = 95% CI) and the standardised mortality rate linear regression (SMR-LR) model (solid red). Both models were fitted to data from 1 January 2014 to 31 December 2019 (indicated by the vertical dashed lines). All calculations use the first quarter of 2021 as the standard population.
@@
Covid-19-attributed deaths were extremely low in 2020 and 2021 (26 in each year). In 2020-23, the number of Covid-19-attributed deaths was close to the central estimate for excess deaths (Figure 3a, blue), accounting for 102% of the estimated excess in 2022 and 82% in 2023, and fell within the 95% CI in both years.
Looking at monthly data (Figure 3b), Covid-19 deaths and excess deaths according to the quasi-Poisson regression model followed similar trends in 2022 and 2023. The number of Covid-19 deaths was below the 95% CI for excess deaths in February and March 2022, and above it from July to September 2022. The unexplained excess in early 2022 could represent unattributed Covid-19 deaths that occurred when the country’s laboratory testing capacity became overwhelmed but before rapid antigen tests became widely available in February [16]. The monthly discrepancies could also partly reflect temporal displacement of mortality due to disruption of seasonal respiratory disease patterns [17].
The number of Covid-19 deaths fell within the 95% CI for the number of excess deaths in each age band (0-59 years, 60-69 years, 70-79 years and 80+ years) and for both sexes in 2022 and 2023 (Figure 4). Note we used these four age bands because data on Covid-19 deaths were available in these bands for comparison (see Supplementary Figure S1 for results disaggregated into finer age groups). The results in Figure 4 mirror similar patterns seen in Australia [18], where there was some positive non-Covid excess in the 70-79-year age band and, despite there being more Covid-19 deaths among men than women, the central estimate for excess deaths was higher for women.
@@
When different pre-pandemic fitting periods were used, the estimated number of excess deaths for 2020-23 according to the quasi-Poisson regression model varied between a low of -1770 [-4973, 924] for a four-year baseline, and a high of 1995 [500, 3368] for a nine-year baseline (Table 1). Expressed as a percentage of expected deaths, these correspond to a low of -1.2% [-3.4%, 0.6%] and a high of 1.4% [0.4%, 2.4%]. The estimated excess according to the standardised mortality rate linear regression model displayed a similar pattern with respect to baseline length. The 95% CI for the quasi-Poisson model contained zero in all cases except for the nine-year baseline.
The Karlinksy and Kobak estimate for 2020-23 of 517 excess deaths [7, 9] falls within the 95% CI for all seven baselines. The Economist estimate of 1026 excess deaths [10] is within the 95% CI for five of the seven baselines, and slightly above it for other two. These estimates are therefore largely consistent with our results. In contrast, Gibson estimated $+4\%$ excess mortality for the three-year period of 2020-22 using a five-year baseline, which translates to approximately 4000 excess deaths [12]. This estimate, which failed to control for changes in the population age distribution, is 4500 deaths higher than our central estimate for this period using a five-year baseline, and 2500 deaths higher than the highest upper limit of the 95% CI for any baseline length between four and 10 years (see Supplementary Table S1).
Gibson’s method is fundamentally flawed because it assumes that expected deaths will be proportional to the total population size. This ignores the fact that, although total population growth slowed during the pandemic, the population aged over 65 years (which is only around 15% of the total population but accounts for 80% of all deaths) continued to grow at a similar rate as before the pandemic (see Supplementary Figure S2).
@@ 4 Discussion
We have estimated the number of excess deaths in Aotearoa New Zealand between January 2020 and December 2023 using two different methods. Importantly, both methods controlled for age, sex and population size (one by including these variables as predictors in a regression model, and one by using an age- and sex-standardised mortality rate). This avoids the potential for systematic errors caused by changes in population size or age structure.
@@
Consistent with other methods [7, 10], both our models estimated negative excess mortality in 2020 and, to a lesser extent, 2021. This was likely due to the effective suppression of Covid-19 (there were only 52 Covid-19 deaths in 2020-21), the elimination of influenza [21], and the reduction in respiratory syncytial virus and other respiratory pathogens during this period [22, 17].
@@
Mortality rates fluctuate from year to year depending on the severity of the winter respiratory illness season and other factors. This means that the fitted baseline for expected mortality can be sensitive to the choice of fitting period. To mitigate this, we ran our model with seven different fitting periods ranging from four to 10 years duration. The limits of the 95% CIs for excess mortality for 2020-23 fell between -3.4% and +2.4% and contained zero for all but one choices of fitting period. Thus our results provide robust evidence that, regardless of the choice of baseline, there was no more than 2.4% cumulative excess mortality between 2020 and 2023 and any net excess that did occur cannot be confidently distinguished from zero.
We have not attempted to make quantitative international comparisons, as these are beyond the scope of this study. Nevertheless, our estimated excess mortality of 204 [-222, 573] deaths per million people for 2020-23 ranks amongst the lowest in the world. Estimated excess mortality for 2020-23 for most OECD countries (with a few notable exceptions including Australia) was in the range 1500–4500 per million [7, 8, 9].
Gibson [12] argued that, by extrapolating raw death counts and not accounting for the abrupt drop in population growth in 2020, the Karlinksy and Kobak method [7] systematically overestimated expected mortality, and therefore underestimated excess deaths. It is true that, where the required data are available, it is preferable to control for total population size and age structure, as these both have a major influence on expected deaths. While Gibson’s method controlled for the former, it did not control for the latter and as a consequence underestimated expected mortality for 2020-22 by approximately 4%. Our results show that, when also controlling for age, the expected number of deaths in the period 2020-23 continued to increase at a similar rate to the pre-pandemic trend. In other words, the ageing population was a more important determinant of the trend in expected deaths than was growth in total population size. This is because the abrupt drop in population growth in 2020 was concentrated in the younger part of the population, which has relatively high levels of migration and low rates of mortality [24]. In contrast, the older part of the population, which has relatively low migration and high mortality rate, continued to grow at a similar rate as in the pre-pandemic period [13]. The main advantage of the method of Karlinsky and Kobak [7] is that it can be applied to a large set of countries to provide a meaningful and useful basis for international comparisons. Despite the systematic errors in application to New Zealand (as for other countries), the Karlinksy and Kobak estimate is within the 95% CIs from our model.
@@
The model lumped all deaths over 95 years old into a single age class. This was necessary because population size data was only available at this level. The size of the over 95 year age group increased from 5130 in 2014-Q1 to 8530 in 2023-Q4. It is likely that the age-distribution within this group became increasingly elderly over time, which would influence the group-specific mortality rate. This is mitigated by the inclusion of an interaction term in the quasi-Poisson model, which allows a group-specific linear trend in the mortality rate. Nevertheless, the model cannot explicitly account for any shift in the age distribution within the group that may have occurred between 2019 and 2023.
@@
Ethics approval was not required for this research as it only uses routinely collected administrative data, which was de-identified and aggregated prior to use.
@@
Data and code to reproduce the analysis are available at https://github.com/michaelplanknz/pandemic-excess-mortality-nz-public. The results shown in the article used raw, unrounded data on death counts. The raw data cannot be published due to privacy concerns relating to small counts. To preserve confidentiality, the repository linked above contains a data set where death counts are randomly rounded (see Supplementary Material Sec. S1.1 for details). Running the code on this dataset will produce results that are similar to those in the article, but not identical and with broader confidence intervals due to the added noise.
@@ Acknowledgements
MJP acknowledges support from Te Niwha Infectious Diseases Research Platform, Institute of Environmental Science and Research, grant number TN/P/24/UoC/MP. The authors are grateful to Jennifer Brown and Philipp Wacker for helpful discussions on confidence intervals for quasi Poisson regression models, and to Rebekah Hennessey and Helen He for assistance with sourcing the deaths data.
@@
Data and Matlab code to reproduce the results in the article are available in the public repository at https://github.com/michaelplanknz/pandemic-excess-mortality-nz-public.
@@
The ERP, at any given point in time, is an estimate based on a starting population and accounting for births and deaths (natural increase/decrease) and net migration [31]. The starting population is the coverage-adjusted resident population established from the census, its associated post-enumeration survey (PES), and an estimate of residents overseas during the census [32].
The starting population is updated after each census, in a process termed “rebasing” [31], with the most recent available being Census 2018. At the time of writing, the Census 2023 based ERP was not yet available, and as with every census update, we expect a revision to the population estimates. For example, in 2018, updating the starting point from Census 2013 to Census 2018 increased the estimated population by 60,000 (approximately 1.2%). The largest difference was among males aged between 30 and 44 years, although 15–29 year olds, and 45–59 year olds also showed notable differences [33].
These differences are driven by a variety of factors, including: (i) inaccuracies in the various census counts and associated coverage adjustments; (ii) conceptual misalignment between what respondents think and the concept of population attempting be measured; and (iii) inaccuracies in the components of population change (births, deaths and migration) between censuses.
@@
$+\beta_{6}(\mathrm{ageCoarse}_{i}\times\mathrm{sex}_{i})+\beta_{7}(\mathrm{ageCoarse}_{i}\times\mathrm{month}_{t})+\log(N_{it})+\log(\mathrm{days}_{t}),$
where $N_{ti}$ is the population size in stratum $i$ at time $t$. The time variable $t$ was an integer-valued variable measured in months since January 2014. The variables $\mathrm{age}_{i}$, $\mathrm{sex}_{i}$ and month of the year ($\mathrm{month}_{t}$) were treated as categorical variables, meaning that a separate coefficient $\beta$ was estimated for each value of the categorical variable (except for the baseline value). This meant estimating 95 coefficients for $\mathrm{age}_{i}$ (one for each age group from 1–2 years up to 95 years and over), 1 coefficient for $\mathrm{sex}_{i}=\mathrm{female}$ and 11 coefficients for $\mathrm{month}_{t}$ (representing February to December).

### v2

@@ Abstract
Key messages
  • Research question: Accounting for New Zealand’s ageing population, how many additional deaths occurred in 2020-23 relative to what would have been expected in the absence of a pandemic?
  • Findings: Excess mortality in New Zealand in 2020-23 was estimated to be 1040, with a 95% confidence interval of [-1134, 2927].
  • Importance: New Zealand experienced one of the lowest rates of pandemic excess mortality in the world.
@@ 1 Introduction
The Covid-19 pandemic caused an estimated 14.8 million deaths globally in 2020 and 2021 [1] Aotearoa New Zealand used a combination of border and community control measures to minimise transmission of SARS-CoV-2 until high vaccine coverage was achieved in late 2021 [2, 3]. Following establishment of the B.1.1.529 (Omicron) variant in January 2022 [4], New Zealand experienced a series of Covid-19 waves [5]. By 31 December 2023, there had been a total of 750 Covid-19-attributed deaths per million people [6], one of the lowest rates in the world.
The number of Covid-19-attributed deaths is one measure of the pandemic’s impact on mortality, but has limitations. It could be influenced by factors such as test availability, how Covid-19-attributed deaths are distinguished from incidental deaths in Covid-19 positive individuals, and processes for recording multiple causes of death. Also, Covid-19-attributed deaths only measure the direct impact of the disease on mortality. This ignores indirect effects, such as deaths that occurred or were prevented as a result of pandemic control measures, and deaths that would have occurred due to another cause within a given time period.
@@
The Karlinsky and Kobak [7] baseline relies on extrapolating the pre-pandemic trend in raw death counts. This may, in some cases, mainly reflect the trend in total population size if the crude mortality rate (i.e. deaths per capita) does not vary greatly. Gibson [12] correctly pointed out that, in such cases, if the trend in total population size abruptly changes, then the Karlinsky and Kobak baseline will be systematically biased. There was a sharp drop in the rate of population growth in New Zealand in 2020 due to reduced international travel. Gibson [12] argued that controlling for population size by fitting to the trend in crude mortality rate, rather than raw deaths, led to a lower baseline than that of Karlinsky and Kobak [7] and, therefore, a higher estimate of excess mortality.
However, the trend in raw deaths may also be influenced by population ageing, which would generally lead to an upward trend in crude mortality rate. For this reason, where age-stratified data are available, it is preferable to control for age when estimating mortality baselines [13]. This will limit international comparisons to jurisdictions where age-stratified data are available, but will provide more accurate estimates. Neither Gibson [12] nor Kung et al. [11] controlled for age. Reliable estimates of pandemic excess mortality in New Zealand that control for the combined effects of population growth and ageing are lacking.
@@
Both methods account for changes in population size and age structure, which means they provide more reliable estimates of excess mortality than methods that do not. We use the quasi-Poisson regression model as our primary method as this allows us to disaggregate results to monthly time periods and by age and sex. We compare aggregated yearly results from this method with the standardised mortality rate linear regression to benchmark our results against a simpler model. We then compare the estimated excess to the number of Covid-19-attributed deaths over time and in different subgroups. This enables us to estimate whether there was any net non-Covid-19 excess mortality during this period.
@@ 2 Methods
We used two methods for constructing a mortality baseline from pre-pandemic data: (1) a generalised linear quasi-Poisson regression model for monthly death counts including age, sex and month of the year as predictors, based closely on the model of the UK’s Office for National Statistics [15]; (2) a linear regression fitted to the age- and sex-standardised annual mortality rate. Quasi-Poisson regression is a natural model for count data whilst allowing for the possibility that the variance may be higher than in a Poisson regression. See Supplementary Material Sec. 1.2 for further details.
@@ 3 Results
Figure 1: Total monthly all-cause deaths in the New Zealand resident population from January 2010 to December 2023 (blue points) along with the quasi-Poisson regression model fitted to data on monthly deaths from January 2014 to December 2019. Vertical dashed lines show the fitting period. Solid blue curve shows the mean monthly deaths according to the fitted model; shaded band shows the 95% CI. Excess mortality is estimated as the difference between actual deaths for 2020-23 and expected deaths for 2020-23 according to the model (mean and 95% CI).
Alt text: Graph showing data for total monthly all-cause deaths in the New Zealand resident population from January 2010 to December 2023, alongside the mean and 95% confidence interval for the quasi-Poisson regression model fitted to the data for the period 2014 to 2019.
In the pre-pandemic period, all-cause monthly deaths in the New Zealand resident population had a pronounced seasonal pattern, with relatively high deaths during the winter (i.e. June–August) respiratory disease season (Figure 1). The quasi-Poisson regression model captured the seasonality and the gradually increasing trend in deaths. The age- and sex-standardised yearly all-cause mortality rate varied between 7.0 and 7.9 per 1000 between 2010 and 2019, with a decreasing trend (Figure 2). The model standardised yearly mortality rate was very similar under the two methods.
Figure 2: Age- and sex-standardised yearly all-cause mortality rate per 1000 people in the New Zealand resident population (open circles), with the results of the quasi-Poisson regression (QPR) model (solid blue curve = mean, shaded blue band = 95% CI) and the standardised mortality rate linear regression (SMR-LR) model (solid red). Both models were fitted to data from 1 January 2014 to 31 December 2019 (indicated by the vertical dashed lines). All calculations use the first quarter of 2021 as the standard population.
Alt text: Graph showing data for the age- and sex-standardised yearly all-cause mortality rate per 1000 people in the New Zealand resident population from 2010 to 2023, alongside the mean and 95% confidence interval for the quasi-Poisson regression model and the standardised mortality rate linear regression model. Both models were fitted to the data for the period 2014 to 2019.
@@
Alt text: First graph shows the yearly number of Covid-19-attributed deaths from 2020 to 2023 alongside the estimated number of excess deaths according to both the quasi-Poisson regression model and the standardised mortality rate linear regression model. Second graph shows the monthly number of Covid-19-attributed deaths from January 2020 to December 2023 alongside the estimated number of excess deaths according the quasi-Poisson regression model.
Covid-19-attributed deaths were extremely low in 2020 and 2021 (26 in each year). In 2022-23, the number of Covid-19-attributed deaths was close to the central estimate for excess deaths (Figure 3a, blue), accounting for 102% of the estimated excess in 2022 and 82% in 2023, and fell within the 95% CI in both years.
Looking at monthly data (Figure 3b), Covid-19 deaths and excess deaths according to the quasi-Poisson regression model followed similar trends in 2022-23. The number of Covid-19 deaths was below the 95% CI for excess deaths in February and March 2022, and above it from July to September 2022. The unexplained excess in early 2022 could represent unattributed Covid-19 deaths that occurred when the country’s laboratory testing capacity became overwhelmed [16]. The monthly discrepancies could also partly reflect temporal displacement of mortality due to disruption of seasonal respiratory disease patterns [17].
The number of Covid-19 deaths fell within the 95% CI for the number of excess deaths in each age band (0-59 years, 60-69 years, 70-79 years and 80+ years) and for both sexes in 2022 and 2023 (Figure 4). Note we used these four age bands because appropriate data on Covid-19 deaths were available for comparison (see Supplementary Figure S1 for results disaggregated into finer age groups). The results in Figure 4 mirror similar patterns seen in Australia [18], where there was some positive non-Covid excess in the 70-79-year age band and, despite there being more Covid-19 deaths among men than women, the central estimate for excess deaths was higher for women.
@@
Alt text: Bar charts showing the yearly number of Covid-19 attributed deaths alongside the estimated number of excess deaths according to the quasi-Poisson regression model. The top two graphs show the numbers of deaths in the 0-59, 60-69, 70-79 and 80+ age groups for 2022 and 2023 respectively. The bottom two graphs show the number of deaths among male and female individuals in 2022 and 2023 respectively.
When different pre-pandemic fitting periods were used, the estimated number of excess deaths for 2020-23 according to the quasi-Poisson regression model varied between a low of -1770 [-4973, 924] for the four-year baseline, and a high of 1995 [500, 3368] for the nine-year baseline (Table S1). Expressed as a percentage of expected deaths, these correspond to a low of -1.2% [-3.4%, 0.6%] and a high of 1.4% [0.4%, 2.4%]. The estimate from our default baseline of six years (2014-19) ranked third out of the seven baselines investigated. The 95% CI for the quasi-Poisson model contained zero in all cases except for the nine-year baseline. The results of the standardised mortality rate linear regression model showed a similar pattern with respect to baseline length. In both models, the variations in estimated excess seen in Table S1 were mainly driven by whether or not the first year in the fitting period had a relatively high all-cause death rate.
The Karlinksy and Kobak estimate for 2020-23 of 517 excess deaths [7, 9] falls within the 95% CI for all seven baselines. The Economist estimate of 1026 excess deaths [10] is very close to our central estimate of 1040 for the six-year baseline, is within the 95% CI for five of the seven baselines, and slightly above it for other two. These estimates are therefore largely consistent with our results. In contrast, Gibson estimated $+4\%$ excess mortality for the three-year period of 2020-22 using a five-year baseline, which translates to approximately 4000 excess deaths [12]. This estimate, which failed to control for changes in population age structure, is 4500 deaths higher than our central estimate for this period using a five-year baseline, and 2500 deaths higher than the highest upper limit of the 95% CI for any baseline length between four and 10 years (see Supplementary Table S1).
@@ 4 Discussion
We have estimated the number of excess deaths in Aotearoa New Zealand between January 2020 and December 2023 using two different methods. Importantly, both methods controlled for age, sex and population size (one by including these variables as predictors in a regression model, and one by using an age- and sex-standardised mortality rate). This avoids the potential for systematic biases caused by changes in population size or age structure.
@@
Consistent with other methods [7, 10], both our models estimated negative excess mortality in 2020 and, to a lesser extent, 2021. This was likely due to the effective suppression of Covid-19 (there were only 52 Covid-19 deaths in 2020-21), the elimination of influenza [21], and the reduction in other respiratory pathogens during this period [22, 17].
@@
Mortality rates fluctuate from year to year depending on the severity of the winter respiratory illness season and other factors. This means that the fitted baseline for expected mortality can be sensitive to the choice of fitting period. To mitigate this, we ran our model with seven different fitting periods ranging from four to 10 years duration. The limits of the 95% CIs for excess mortality for 2020-23 fell between -3.4% and +2.4% and contained zero for all but one choice of fitting period. Thus our results provide robust evidence that, regardless of the choice of baseline, there was no more than 2.4% cumulative excess mortality between 2020 and 2023 and any net excess that did occur cannot be confidently distinguished from zero.
We have not attempted to apply our method to estimate excess mortality in countries other than New Zealand, as this is beyond the scope of the study. Nevertheless, our estimated excess mortality of 204 [-222, 573] deaths per million people for 2020-23 ranks amongst the lowest in the world. Estimated excess mortality for 2020-23 for most OECD countries (with a few notable exceptions including Australia) was in the range 1500–4500 per million [7, 8, 9].
Gibson [12] argued that, by extrapolating raw death counts and not accounting for the abrupt drop in population growth in 2020, the Karlinksy and Kobak method [7] systematically overestimated expected mortality, and therefore underestimated excess deaths. However, although population growth slowed dramatically during the pandemic, this was largely a result of reduced migration, which occurs predominantly in younger groups [24]. The population over 65 years of age (which is only around 15% of the total population but accounts for 80% of deaths) continued to grow at a rate similar to before the pandemic (see Supplementary Figure S2). By controlling for total population size but not for age, Gibson’s method introduced a systematic bias and consequently underestimated expected mortality for 2020-22 by approximately 4%. The main advantage of the method of Karlinsky and Kobak [7] is that it can be applied to a large set of countries to provide a meaningful and useful basis for international comparisons. Despite not controlling for population size or age, the Karlinksy and Kobak estimate was within the 95% CIs from our model. This is because the pre-pandemic trend in death counts was driven more by population ageing (which continued throughout the pandemic) than by population growth (which did not).
@@
Our model lumped all deaths over 95 years old into a single age class. This was necessary because population size data was only available at this level. The size of the over 95 year age group increased from 5130 in 2014-Q1 to 8530 in 2023-Q4. It is likely that the age-distribution within this group became increasingly elderly over time, which would influence the group-specific mortality rate. This is mitigated by the inclusion of an interaction term in the quasi-Poisson model, which allows a group-specific linear trend in the mortality rate. Nevertheless, the model cannot explicitly account for any shift in the age distribution within the group that may have occurred between 2019 and 2023.
@@
Ethics approval was not required for this research as it only used routinely collected administrative data, which was de-identified and aggregated prior to use.
@@
Data and code to reproduce the analysis are available at https://dx.doi.org/10.5281/zenodo.15107131. The results shown in the article used raw, unrounded data on death counts. The raw data cannot be published due to privacy concerns relating to small counts. To preserve confidentiality, the repository linked above contains a data set where death counts were randomly rounded (see Supplementary Material Sec. S1.1 for details). Running the code on this dataset will produce results that are similar to those in the article, but not identical and with broader confidence intervals due to the added noise.
@@ Acknowledgements
MJP acknowledges support from Te Niwha Infectious Diseases Research Platform, Institute of Environmental Science and Research, grant number TN/P/24/UoC/MP. The authors are grateful to Jennifer Brown and Philipp Wacker for helpful discussions on confidence intervals for quasi Poisson regression models, to Rebekah Hennessey and Helen He for assistance with sourcing the deaths data, and to two anonymous peer reviewers for helpful comments on an earlier version of this manuscript.
@@
Data and Matlab code to reproduce the results in the article are archived and publicly available https://dx.doi.org/10.5281/zenodo.15107131.
@@
The ERP, at any given point in time, is an estimate based on a starting “base population” and accounting for subsequently observed births, deaths (natural increase/decrease) and net migration [31]. This base population is the coverage-adjusted resident population established from the most recent census, its associated post-enumeration survey (PES), and an estimate of residents overseas during the census [32].
The base population is updated after each census, in a process termed “rebasing” [31], with the most recent available being Census 2018. At the time of writing, the Census 2023 based ERP was not yet available. As with every census update, we expect a revision to the population estimates for 2023 (and to a lesser extent 2019-2022) when the new census becomes available. For example, updating from Census 2013 to Census 2018 increased the estimated population for 2018 by 60,000 (approximately 1.2%). The largest difference was among males aged between 30 and 44 years, although 15–29 year olds, and 45–59 year olds also showed notable differences [33]. These differences were driven primarily by inaccuracies in migration estimates and, to a lesser extent, births and deaths between censuses. They may also have been influenced by inaccuracies in the various census counts and associated coverage adjustments; and potential mismatch between the definition of a New Zealand resident and census respondents’ interpretation of this.
@@
$+\beta_{6}(\mathrm{ageCoarse}_{i}\times\mathrm{sex}_{i})+\beta_{7}(\mathrm{ageCoarse}_{i}\times\mathrm{month}_{t})$
$+\log(N_{it})+\log(\mathrm{days}_{t}),$
where $N_{it}$ is the population size in stratum $i$ at time $t$. The time variable $t$ was an integer-valued variable measured in months since January 2014. The variables $\mathrm{age}_{i}$, $\mathrm{sex}_{i}$ and month of the year ($\mathrm{month}_{t}$) were treated as categorical variables, meaning that a separate coefficient $\beta$ was estimated for each value of the categorical variable (except for the baseline value). This meant estimating 95 coefficients for $\mathrm{age}_{i}$ (one for each age group from 1–2 years up to 95 years and over), 1 coefficient for $\mathrm{sex}_{i}=\mathrm{female}$ and 11 coefficients for $\mathrm{month}_{t}$ (representing February to December).